OpenStudy (anonymous):

Standard form? Is this right? 37. slope = ; (2, 8) y-8=3/4(x-2) y-8=3/4x-6/4 y=3/4x+26/4 4y=3x+26 -3x+4y=26

OpenStudy (anonymous):

hey have you tried gegebra

OpenStudy (anonymous):

sorry geogebra

OpenStudy (anonymous):

It can do standard form? I've used it for graphing but this isn't about graphing.

OpenStudy (anonymous):

it can help with slope. Let me look at your problem and see if I can see if you are right. ahh gotcha you are wanting to see if you wrote it down right

OpenStudy (anonymous):

Yeah I wasn't sure if I made it into standard form correctly given slope and one point.

OpenStudy (anonymous):

it looks right but not positive

OpenStudy (anonymous):

Ok well then I guess I'll keep waiting to see if someone can tell me if it is or not. It's just for a study guide but it is worth a small part of my grade and I'd really like to do good on it.

OpenStudy (phi):

I don't see what they gave you for the slope. But if the slope is ¾ , it looks ok, but often standard form means: variables in alphabetical order and leading coefficient on the first variable is POSITIVE. To be safe, I would write -3x+4y=26 as 3x - 4y = - 26

OpenStudy (anonymous):

Oh right! It's been a long time since I've done these and yes slope was 3/4!

OpenStudy (anonymous):

If I do a few more could you tell me if they're right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

slope = -2/5 ; (3, -9) y+9=-2/5(x-3) y+9=-2/5x+6/5 This is where I get stuck. Do I multiply the whole thing by 5 or do I subtract 9 on both sides first?

OpenStudy (phi):

You can take either step. But standard form does not use fractions, and at some point we will have to multiply by 5 to "clear the denominator" In fact, I would make that the very first step for this problem. \[ y+9 = -\frac{2}{5}(x-3) \\ 5y +45 = -2(x-3)\] but if we pick up with \[ y+9=-\frac{2}{5}x+\frac{6}{5} \] I would multiply by 5 to get \[ 5y +45= -2x + 6\]

OpenStudy (anonymous):

So my answer would be 2x+5y=-39

OpenStudy (phi):

yes

OpenStudy (anonymous):

Awesome thank you I think I got it now!

OpenStudy (phi):

notice for do I subtract 9 on both sides first? \[ y+9=-\frac{2}{5}x+\frac{6}{5} \\ y = -\frac{2}{5}x+\frac{6}{5} -9 \] again we could multiply by 5. or we can add 2/5 x to both sides to get \[ \frac{2}{5}x + y = \frac{6}{5}-9 \] we could simplify 6/5 - 9 to -39/5. or we could multiply by 5 (on both sides) \[ 2x + 5y = 6-45 \\ 2x + 5y = -39 \]